A Las Vegas roulette wheel has 38 numbered slots. Roulette wheels are calibrated such that each outcome is

equally likely. To play roulette, you choose a number; you win$35 if the ball ends up in that slot and lose $1 otherwise.

d). Estimate the chance that you will win more than $12 in 100 plays.
Answer #1

The probability of winning is

p = P(win) = 1/38

Let X is a random variable shows the winning amount. Here X can take value 35 and -1. So,

P(X=35) = P(win) = 1/38

P(X=-1) = 1-P(win) = 37/ 38

The expected value of X is

E(X) = (1/38) * 35 + (37/38) * (-1)= -2 / 38 =-0.0526

Now,

1188 37 . 352 38 38 E(X2) (1)2 38

The standard deviation is

1188 4 5.5911 g E(X2) E(X2 382 38

d)

Let T shows the total amount you win in 100 plays. Using CLT, T will have normal distribution with mean and SD as follow,

100 E(X 100(0.0526) -5.26 Дт

aT 100 o 10.(5.5911) 55.911

The z-score for T = 12 is

12 5.26 T 55.911 0.31

The  chance that you will win more than $12 in 100 plays is

P(T> 12)= P(z> 0.31)1- P(z<0.31) 1-0.6217 0.3783